OpenStudy (anonymous):

I WILL FAN AND MEDAL! A recipe that makes 6 servings calls for 11/5 cups of juice. How many cups of juice are there per serving?

OpenStudy (jdoe0001):

\(\large \begin{array}{ccllll} servings&cups \\\hline\\ 6&\frac{11}{5}\\ 1&x \end{array}\implies \cfrac{6}{1}=\cfrac{\frac{11}{5}}{x}\) solve for "x"

OpenStudy (jdoe0001):

recall that \(\large \cfrac{\quad \frac{a}{b}\quad }{\frac{c}{d}}\implies \cfrac{a}{b}\cdot \cfrac{d}{c}\)

OpenStudy (anwaarullah):

11/5*6=11/30=.366